Consider pepper jack cheese and parmesan cheese markets for the US. The domestic market equilibrium price and quantity for pepper jack is $10 and 100 respectively. On the other hand The domestic market equilibrium price and quantity for parmesan is $15 and 150 respectively. The world price of pepper iack is $12 and world price of parmesan is $10. Assume that the US does not have any power to effect the price level in the world. Assume the US opens up to international production for both of these goods. Answer the following 3 questions according to this passage. Which of the following is true? US is the exporter of pepper jack and importer of parmesan. O We dont have enough information to answer this question US is the exporter of parmesan and importer of pepper jack. QUESTION 11 Now assume that quantity supplied for pepper jack is 120 and quantity demanded for pepper jack is 90 under world price. What are the gains from trade for the US in pepper jack market? O 40 10 30 20
Answer #1

If a country has comparative advantage in production of a good then the country has lower opportunity cost of producing that good. Nations that produce according to their comparative advantage are maximizing the benefits that they receive from trade and consequently, their national welfare. The benefit that one country accrues from trade is called gains from trade.

According to the data given, the domestic market price for paper jack is $10 while the international price is $12. then domestic producers have comaprative advantage in paper jack. Hence, US will export paper jack and will earn $2 profit per unit.

On the other hand, the domestic price of parmesan is $15 and the world price is $12. Then the country can gain $3 per unit if it buys parmesan fromthe world. Then US will import parmesan.

Therefore, the correct option is

  • exporter of paperjack-importer of parmesan

====================================================================

The gains from trade in case of export is given by the area ABC below

ABC=\frac{1}{2}\times (12-10)\times (120-90)=30

Price Supply 12 World Supply Demand 0 90 120 Quantity

Therefore, the correct option is

  • 30